Let individual i have the utility function Ui(x,y) = x^1/2 * y Suppose the initial endowment is x = 16 and y = 2. a. Determine 4 combinations (x,y) for which Ui=8 ; Ui=16. b. What is the minimum amount of x that i would have to be compensated to introduce him to give up 1 unit of y?
Answer #1

The utility function is given as U; = 1/2 .

(a) For utility level be 8, we have 8 = 1/2 or 64 = ry or x = \frac{64}{y^2} . For y=1, x = \frac{64}{1^2} = 64 . For y=2, x = \frac{64}{2^2} = 16 . For y=4, x = \frac{64}{4^2} = 4 . For y=8, x = \frac{64}{8^2} = 1 .

Hence, the required four combinations of (x,y) for U=8 is fr, ):{{64, 1), (16, 2), (4, 4), (1,8)} .

For utility level be 16, we have 16 = 1/2 or 256 = ry or x = \frac{256}{y^2} . For y=1, x = \frac{256}{1^2} = 256 . For y=2, x = \frac{256}{2^2} = 64 . For y=4, 91 = - . For y=8, f %3D- 9 . Additionally, for y=16, x = \frac{256}{16^2} = 1 .

Hence, the required four combinations of (x,y) for U=16 is (x,y) : \left \{ (256,1) , (64,2) , (16,4) , (4,8) \right \} .

(b) For the given endowment, we have U_i = 16^{1/2}*2 = 8 . From the set of bundles fr, ):{{64, 1), (16, 2), (4, 4), (1,8)} , we can see that, consumer is endowed at second point. To go to the first point, thereby reducing the consumption of y by 1 unit, the consumer must be compensated with 64 minus 16, ie 48 units of x. Hence, the minimum amount of x is 48 units, which have to be compensated to induce consumer to give up 1 unit of y.